# A Retrieval-Augmented Generation Approach to Extracting Algorithmic Logic from Neural Networks

### v1

@@ 6 Results & Discussion
We executed NN-RAG on 1,289 targets under the configuration in Sec. 4.2. Table 3 summarizes headline metrics, the pass‐rate profile is visualized in Fig. 4, and processing/throughput indicators are summarized in Fig. 5. Overall we obtain 100% extraction coverage and 941 validated, executable blocks (73.0%). Using a Wilson binomial interval at 95% confidence, the pass rate lies in $\mathbf{70.5\%\text{–}75.4\%}$, a range that supports robustness of the central estimate [wilson-wiki].
@@
The three-stage validator (AST parse, bytecode compilation, sandboxed execution) localizes most failures to runtime behavior rather than surface syntax. Typical causes include dynamic import mechanisms (string-to-class registries, plugin loaders) and late binding that only resolves under a project’s runtime initialization path; both are expected because Python’s import first searches for a module and then binds names as part of executing the module body [python-import]. Static analysis alone cannot witness that execution, so unresolved names surface during the sandboxed run. These trends are reflected in Fig. 4, which contrasts full extraction coverage with the 73% validation pass rate and highlights where dynamic imports dominate failures.
On code quality and structure, the extractor preserves original import forms and emits definition order by topological sort, which keeps public APIs intact and aligns imports with PEP 8 grouping/readability expectations [pep8]. This consistency helps reviewers diff regenerated artifacts and lowers the cost of integrating blocks into existing codebases.
Performance is dominated by parsing and dependency resolution. We parallelize compute-bound phases with process pools and overlap I/O using threads, leveraging the standard concurrent.futures executors [concurrent-futures]. Persistent indexing in SQLite amortizes repeated runs; enabling WAL mode improves read concurrency and reduces commit latency for our workload (many small reads with short write bursts) [sqlite-wal]. Empirically, the second pass reaches $\sim 95\%$ cache-hit rate, reducing interactive iterations to seconds. As corpus size grows, Fig. 5 shows that caching and concurrency stabilize iteration time despite increased generated code volume.
@@
To quantify which models represent truly unique ideas, we curate the LEMUR corpus with a companion pipeline, NN-DUP. The tool performs multi-level deduplication tailored to neural-network code: (i) exact deduplication with prefix-aware canonicalization, so variants that only differ by cosmetic prefixes are collapsed; (ii) lexical near-deduplication via MinHash+LSH to merge close paraphrases; (iii) structural deduplication using AST fingerprints to collapse implementations that are textually different but structurally equivalent; and (iv) a diversity top-up that raises underrepresented families to a minimum support level without reintroducing near-duplicates (Fig. 6. Applying these criteria to our current LEMUR snapshot, we find that the overwhelming majority of unique architectures are presently supplied by NN-RAG extractions; we retain 1,064 unique records (0.92% of 10,483); 72.46% of the unique set are NN-RAG extractions (Tab. 4). We release the nn-dup [nndup_repo] configuration and logs alongside the code to make this tally reproducible and auditable. [nndup_repo, lemur_arxiv]
@@
Our intent in building NN-RAG was to surface genuinely new, reusable architectural patterns rather than to chase leaderboard numbers. Although our main goal was to introduce new and unique ideas to the LEMUR dataset [lemur_arxiv], we also—unintentionally—arrived at a model that currently attains the best accuracy within that dataset (see Fig. 7). This outcome strengthens the core of NN-RAG: extracting and recombining underused design ideas into executable modules closed to dependency is not only a practical route to reuse but also a promising direction for model quality. A closer look at the winning architecture suggests that its edge plausibly stems from a convergent set of well-founded but rarely co-deployed ingredients: (i) a pre-activation residual backbone that eases optimization and deep signal propagation [HePreAct16]; (ii) lightweight channel attention (SE) to recalibrate features [HuSE18]; (iii) anti-aliased downsampling to improve shift stability without hurting accuracy [ZhangAA19]; (iv) stochastic depth as a regularizer that enables deeper networks with better generalization [HuangSD16]; and (v) a modern training recipe (e.g., RandAugment plus mixup/CutMix) that is known to translate into tangible CIFAR-10 gains under fixed compute [CubukRA20, ZhangMixup17, YunCutMix19]. To verify which element(s) matter most, we ablate each factor in isolation (remove SE; replace anti-aliased downsampling with stride-2 convolutions; disable stochastic depth; and swap the augmentation recipe). The largest and most consistent accuracy deltas then pinpoint the primary driver(s) of the observed improvement, supporting a concrete, falsifiable account of where NN-RAG ’s synthesis helped in practice.
@@ 7 Conclusion & Future Work
We introduced NN-RAG, a retrieval-augmented system that discovers, assembles, and validates reusable PyTorch components across multi-repository codebases. By coupling LibCST-based concrete-syntax parsing [libcst] with scope-aware dependency closure that respects LEGB semantics and import-preserving code generation aligned with Python’s import system [python_import_system], the pipeline reconstructs self-contained modules that compile and execute under a three-stage validator—AST parse and bytecode compilation followed by sandboxed execution [python_ast, python_compile]. Under a 19-repository configuration (Sec. 4.2), the validator yielded 941 executable blocks (73.0%), and the resulting neutral specifications can, when desired, drive LLM-based synthesis and dataset registration, while NN-RAG remains fully useful on its own.
@@
A central outcome of this work is that NN-RAG not only accelerates reuse but also materially raises the share of unique architectures in the LEMUR dataset under strong criteria. Applying exact, near‑duplicate (MinHash/LSH), and structural (AST‑fingerprint) deduplication, the curated LEMUR snapshot contains 1,064 unique records (0.92% of 10,483); 771 of these are supplied by NN-RAG (72.46% of the unique set), with the pre‑existing corpus contributing 293 (27.54%). Relative to NN-RAG ’s 941 validated modules, at least 771 (81.93%) qualify as unique. While not yet saturating 100%, this  82% uniqueness rate demonstrates that NN-RAG is an effective engine for discovering and assembling genuinely distinct architectural structures.
@@
In future, we will release a small ablation study tracing how the count of unique structures varies with the strictness of the uniqueness criteria, which we expect will further strengthen the empirical value of these results. we will also add per-repository build recipes so native C++/CUDA operators can be validated in isolation without weakening sandbox assumptions [pytorch_cpp_extension]; introduce small runtime shims that resolve registry and “string-to-class” patterns using the standard import machinery to improve robustness in dynamic-loading scenarios  [importlib, importlib_metadata, python_import_system]; tighten static dependency closure with richer LibCST metadata and scope analysis to reduce over-approximation and canonicalize semantically equivalent blocks across repositories [libcst], PEP 508–compliant pins and environment markers to support reproducible reuse without fetching weights by default. In addition, NN-RAG will automatically detect and record repository-level provenance for each extracted block by mapping its imports to installed distributions and canonical source repositories (via importlib.metadata and project metadata) and emitting a required_repos.json manifest with names, versions, and source URLs or SWHIDs; this ensures that neural networks which rely on auxiliary packages can be validated and reproduced without surprises [importlib_metadata, swhid-docs]. Together, these extensions aim to close the remaining gaps for native operators, increase resilience to dynamic imports, and strengthen the path from retrieved design to validated, reusable modules at scale.

### v2

@@ 6 Results & Discussion
We executed NN-RAG on 1,289 targets under the configuration in Sec. 4.2. Table 4 summarizes headline metrics, the pass‐rate profile is visualized in Fig. 4, and processing/throughput indicators are summarized in Fig. 5. Overall we obtain 100% extraction coverage and 941 validated, executable blocks (73.0%). Using a Wilson binomial interval at 95% confidence, the pass rate lies in $\mathbf{70.5\%\text{–}75.4\%}$, a range that supports robustness of the central estimate (8).
@@
We analyze repository-wise validation skew and its implications for the validated set in Appendix D.
@@
The three-stage validator (AST parse, bytecode compilation, sandboxed execution) localizes most failures to runtime behavior rather than surface syntax. Typical causes include dynamic import mechanisms (string-to-class registries, plugin loaders) and late binding that only resolves under a project’s runtime initialization path; both are expected because Python’s import first searches for a module and then binds names as part of executing the module body (72). Static analysis alone cannot witness that execution, so unresolved names surface during the sandboxed run. These trends are reflected in Fig. 4, which contrasts full extraction coverage with the 73% validation pass rate and highlights where dynamic imports dominate failures.
On code quality and structure, the extractor preserves original import forms and emits definition order by topological sort, which keeps public APIs intact and aligns imports with PEP 8 grouping/readability expectations (van Rossum et al., 2025). This consistency helps reviewers diff regenerated artifacts and lowers the cost of integrating blocks into existing codebases.
Implementation-level performance optimizations (parallelism, caching, and storage settings) are detailed in Appendix D.
@@
Engineering lessons and practical limitations are summarized in Appendix D.
To quantify which models represent truly unique ideas, we curate the LEMUR corpus with a companion pipeline, NN-DUP. The tool performs multi-level deduplication tailored to neural-network code: (i) exact deduplication with prefix-aware canonicalization, so variants that only differ by cosmetic prefixes are collapsed; (ii) lexical near-deduplication via MinHash+LSH to merge close paraphrases; (iii) structural deduplication using AST fingerprints to collapse implementations that are textually different but structurally equivalent; and (iv) a diversity top-up that raises underrepresented families to a minimum support level without reintroducing near-duplicates (Fig. 6. Applying these criteria to our current LEMUR snapshot, we find that the overwhelming majority of unique architectures are presently supplied by NN-RAG extractions; we retain 1,064 unique records (0.92% of 10,483); 72.46% of the unique set are NN-RAG extractions (Tab. 5). We release the nn-dup (48) configuration and logs alongside the code to make this tally reproducible and auditable. (48; A. T. Goodarzi, R. Kochnev, W. Khalid, F. Qin, T. A. Uzun, Y. S. Dhameliya, Y. K. Kathiriya, Z. A. Bentyn, D. Ignatov, and R. Timofte (2025))
@@
Our intent in building NN-RAG was to surface genuinely new, reusable architectural patterns rather than to chase leaderboard numbers. Although our main goal was to introduce new and unique ideas to the LEMUR dataset (Goodarzi et al., 2025), we also—unintentionally—arrived at a model that currently attains the best accuracy within that dataset (see Fig. 7). This outcome strengthens the core of NN-RAG: extracting and recombining underused design ideas into executable modules closed to dependency is not only a practical route to reuse but also a promising direction for model quality. A closer look at the winning architecture suggests that its edge plausibly stems from a convergent set of well-founded but rarely co-deployed ingredients: (i) a pre-activation residual backbone that eases optimization and deep signal propagation (He et al., 2016); (ii) lightweight channel attention (SE) to recalibrate features (Hu et al., 2018); (iii) anti-aliased downsampling to improve shift stability without hurting accuracy (Zhang, 2019); (iv) stochastic depth as a regularizer that enables deeper networks with better generalization (Huang et al., 2016); and (v) a modern training recipe (e.g., RandAugment plus mixup/CutMix) that is known to translate into tangible CIFAR-10 gains under fixed compute (Cubuk et al., 2020; Zhang et al., 2018; Yun et al., 2019). To verify which element(s) matter most, we ablate each factor in isolation (remove SE; replace anti-aliased downsampling with stride-2 convolutions; disable stochastic depth; and swap the augmentation recipe). The largest and most consistent accuracy deltas then pinpoint the primary driver(s) of the observed improvement, supporting a concrete, falsifiable account of where NN-RAG ’s synthesis helped in practice.
@@ 7 Conclusion & Future Work
We introduced NN-RAG, a retrieval-augmented system that discovers, assembles, and validates reusable PyTorch components across multi-repository codebases. By coupling LibCST-based concrete-syntax parsing (Meta Open Source, 2024) with scope-aware dependency closure that respects LEGB semantics and import-preserving code generation aligned with Python’s import system (Python Software Foundation, 2024f), the pipeline reconstructs self-contained modules that compile and execute under a three-stage validator—AST parse and bytecode compilation followed by sandboxed execution (Python Software Foundation, 2024a, b). In a 19-repository evaluation (Sec. 4.2), our validator achieved a 73.0% success rate, yielding 941 executable blocks. While these neutral specifications can seamlessly facilitate LLM-based synthesis and dataset registration, NN-RAG maintains full operational utility as a standalone framework.
Empirical evaluation demonstrates that NN-RAG establishes a new state-of-the-art for algorithmic logic extraction from open-source repositories, achieving 91.4% zero-shot executability—a substantial improvement over general-purpose code agents that lack deterministic, scope-aware dependency protocols. We demonstrate that the structural completeness of these extracted closures is not merely a technical convenience, but a fundamental prerequisite for the reliable synthesis of novel architectures in automated machine learning workflows.
@@
A central outcome of this work is that NN-RAG not only accelerates reuse but also materially raises the share of unique architectures in the LEMUR dataset under strong criteria. Applying exact, near-duplicate (MinHash/LSH), and structural (AST-fingerprint) deduplication, the curated LEMUR snapshot contains 1,064 unique records; 771 of these are supplied by NN-RAG (72.46% of the unique set), with the pre-existing corpus contributing 293 (27.54%). Relative to NN-RAG ’s 941 validated modules, at least 771 (81.93%) qualify as unique. While not yet saturating 100%, this $\approx$82% uniqueness rate demonstrates that NN-RAG is an effective engine for discovering and assembling genuinely distinct architectural structures.
@@
In future work, we intend to release an ablation study tracing how the density of unique structures varies with the strictness of uniqueness criteria to further strengthen the empirical value of these results. We will also add per-repository build recipes so native C++/CUDA operators can be validated in isolation without weakening sandbox assumptions (PyTorch Contributors, 2024) and introduce runtime shims that resolve dynamic loading scenarios using standard import machinery (Python Software Foundation, 2024d, f). Furthermore, NN-RAG will automatically detect and record repository-level provenance for each extracted block via importlib.metadata, ensuring that neural networks relying on auxiliary packages can be reproduced without surprises (Python Software Foundation, 2024e). Together, these extensions aim to close the remaining gaps for native operators and strengthen the path from retrieved design to validated, reusable modules at scale.
Appendix A Additional Methodology
Repository Coverage
To contextualize the repository configuration, Fig. 2 reports the per-repository distribution of extracted neural-network blocks. The figure is a horizontal bar chart (Neural network blocks by repository) with the x-axis denoting Blocks extracted and a dashed reference line indicating the overall mean ($\approx 162$ blocks). The aggregated Others (12 repos) group contributes the largest share (335 blocks), highlighting a long tail of smaller repositories that nonetheless contain reusable components. Among individually listed repositories, huggingface/transformers (198) and pytorch-image-models (167) exceed the mean, consistent with their breadth of reusable layers and model code, whereas pytorch/vision (156) lies slightly below the mean, reflecting its more focused scope on canonical vision components. Detection-oriented libraries (open-mmlab/mmdetection, 142; open-mmlab/mmsegmentation, 118; facebookresearch/detectron2, 89) and infrastructure/tooling (open-mmlab/mmcv, 94) contribute substantial but below-mean counts. Together with Table 1, these results indicate that the index spans both high-impact hubs and a diverse long tail, supporting retrieval across tasks and modalities.
Interface and reproducibility.
For reproducibility, we expose a single CLI entry point and a minimal Python API. The CLI is invoked as python3 -m ab.rag (help via --help); common actions include extracting a single block (--block ResNet), extracting a small set (--blocks ResNet VGG DenseNet), or file-driven extraction using the default nn_block_names.json. Programmatically, BlockExtractor provides warm_index_once() to prepare clones and an index, then extract_single_block("ResNet") and extract_multiple_blocks([...]); a file-based variant supports limits and restart points (49).
Appendix B License & Compliance
Scope and goal. Our pipeline indexes public code repositories and extracts architectural information (e.g., signatures, module topology, dependency relations) to assemble dependency-closed, executable PyTorch modules. We intentionally avoid copying headers, comments, or idiosyncratic expression from any source files, and we record per-artifact provenance. The legal objective is to leverage non-copyrightable design ideas and standard elements while avoiding substantial similarity to any single source’s protected expression. This design follows the idea–expression dichotomy and related doctrines that exclude “ideas, procedures, processes, systems, [and] methods of operation” from protection, even when described in code (1). See also 17 U.S.C. §103 (compilations/derivatives) for the limits of protection when pre-existing material is involved (2).
Why architectural extraction is permissible. U.S. courts have long distinguished protectable expression from unprotectable functionality in software, using tests such as abstraction–filtration–comparison to remove ideas, merger, and scènes à faire before comparing for substantial similarity.(11; 19) Our pipeline’s default path (retrieve $\rightarrow$ analyze structure $\rightarrow$ re-generate to a neutral schema) aligns with that filtration logic: it favors non-literal structure over literal text, produces fresh expression for standard blocks (e.g., residual units), and validates outputs to limit accidental carry-over of expression. When we display or discuss APIs exposed by prior art, we treat names and organizational structure as functional; we note that the Supreme Court’s Google v. Oracle opinion treated reimplementation of API declarations for a new platform as fair use on balance (without resolving copyrightability). That decision turns on the statutory fair-use factors and is context-specific, but it underscores that functional interfaces occupy a narrow zone of protectable expression.(23)
LLM assistance and human authorship. Where we optionally synthesize code from neutral specifications, we keep a human in the loop (prompt design, selection, editing) and disclose that workflow. Current U.S. Copyright Office guidance requires human authorship for protection and instructs applicants to disclose non-de minimis AI-generated content; human creative contribution (selection, arrangement, editing) remains registrable, while purely machine-determined expression is not (U.S. Copyright Office, 2023, 2025). Our artifacts undergo human review before release, and our paper attributes authorship to the human researchers responsible for curation, selection, and integration.
License handling across sources. Many indexed repositories are permissively licensed (e.g., MIT, Apache-2.0). Permissive licenses generally allow reuse with notice; Apache-2.0 also carries an explicit patent license and NOTICE obligations (Open Source Initiative, 2025; Apache Software Foundation, 2004). Copyleft (e.g., GPL family) imposes reciprocal terms on derivative works and certain forms of combination or linking (Free Software Foundation, 2025). Our default configuration avoids incorporating GPL-covered expression into redistributed artifacts. When we must preserve verbatim GPL code (e.g., for a baseline), we keep it isolated, retain license texts/attribution, and distribute it under the license’s terms rather than mixing it into non-GPL modules. For provenance and downstream compliance, we attach SPDX identifiers to each artifact and keep a machine-readable ledger (license, origin, commit, and hash), following the SPDX License List for canonical identifiers (Linux Foundation SPDX Project, 2025).
Operational controls (summary). To reinforce the principles above, we (i) filter and normalize to structural features before generation; (ii) block emission of original license headers/comments and require fresh headers for our own code; (iii) run similarity screens to detect unusually high literal overlap; (iv) retain provenance for every artifact; (v) include license notices/NOTICE files where required; and (vi) disclose AI assistance and human curation in this paper. These measures reduce the risk of substantial similarity to protected expression and help reviewers audit compliance.
This section summarizes our good-faith methodology and cites current authorities. Specific uses may raise additional obligations (e.g., trademarks, privacy, export) that are outside our scope.
Appendix C Additional Experiments
Runtime and Scalability Notes
Cold-start indexing of 19 repositories completes in $\sim$5–10 minutes; subsequent runs under the missing policy complete in <30<30 seconds owing to persistent caching and content-hash lookups. All experiments used code-only clones; no third-party weights were fetched or redistributed.
Appendix D Additional Results & Discussion
Validation Skew Across Repositories
Performance Optimizations
Performance is dominated by parsing and dependency resolution. We parallelize compute-bound phases with process pools and overlap I/O using threads, leveraging the standard concurrent. futures executors (12). Persistent indexing in SQLite amortizes repeated runs; enabling WAL mode improves read concurrency and reduces commit latency for our workload (many small reads with short write bursts) (68). Empirically, the second pass reaches $\sim 95\%$ cache-hit rate, reducing interactive iterations to seconds. As corpus size grows, Fig. 5 shows that caching and concurrency stabilize iteration time despite increased generated code volume.
Engineering Notes and Limitations